A population consists of the five measurements 2, 6, 8, 0, and 1.



1. How many different samples of size 3 can be drawn from the population (no replacement)?

2. Construct the sampling distribution of the sample means.

3. What is the mean of the sampling distribution of the sample means UX?

4. What is the standard deviation of the sampling distribution of the sample means?


1
Expert's answer
2021-06-03T03:53:46-0400

 1.

We have population values "2,6,8,0,1" population size "N=5" and sample size "n=3."Thus, the number of possible samples which can be drawn without replacement is


"\dbinom{N}{n}=\dbinom{5}{3}=10"


2.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 2,6, 8 & 16/3 \\
 \hdashline
 2 & 2,6,0 & 8/3 \\
 \hdashline
 3 & 2,6,1 & 9/3 \\
 \hdashline
 4 & 2,8,0 & 10/3\\
 \hdashline
 5 & 2,8,1 & 11/3 \\
\hdashline
 6 & 2,1,0 & 3/3 \\
 \hline
7 & 6,8,1 & 15/3 \\
 \hline
 8 & 6,8,0 & 14/3 \\
 \hline
9 & 6,1,0 & 7/3 \\
 \hline
10 & 8,1,0 & 9/3 \\
 \hline
 \hline
\end{array}"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 3/3 & 1& 1/10 & 3/30 & 9/90 \\
 \hdashline
 7/3 & 1 & 1/10 & 7/30 & 49/90 \\
 \hdashline
 8/3 & 1 & 1/10 & 8/30 & 64/90 \\
 \hdashline
 9/3 & 2 & 2/10 & 18/30 & 162/90 \\
 \hdashline
 10/3 & 1 & 1/10 & 10/30 & 100/90 \\
 \hdashline
 11/3 & 1& 1/10 & 11/30 & 121/90 \\
 \hdashline
 14/3 & 1& 1/10 & 14/30 & 196/90 \\
 \hdashline
 15/3 & 1& 1/10 & 15/30 & 225/90 \\
 \hdashline
 16/3 & 1& 1/10 & 16/30 & 256/90 \\
 \hdashline
 Total & 10 & 1 & 102/30 & 1182/90 \\ \hline
\end{array}"



3.

Mean


"\mu=\dfrac{2+6+8+0+1}{5}=3.4"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{102}{30}=3.4"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.



"E(\bar{X})=3.4=\mu"


4.

Variance


"\sigma^2=\dfrac{1}{5}\big((2-3.4)^2+(6-2.4)^2+(8-3.4)^2""(0-3.4)^2+(1-3.4)^2\big)=9.44"


Standard deviation



"\sigma=\sqrt{\sigma^2}=\sqrt{9.44}\approx3.0725"




"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"




"=\dfrac{1182}{90}-(\dfrac{102}{30})^2=\dfrac{118}{75}\approx1.573333"




"\sqrt{Var(\bar{X})}=\sqrt{\dfrac{118}{75}}\approx1.254326"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{9.44}{3}(\dfrac{5-3}{5-1})""=\dfrac{9.44}{6}=\dfrac{118}{75}\approx1.573333, True"